OpenStudy (anonymous):

2inch squares are cut from each corner of a square peice of sheet metal and then folded up to create an open top box. if the volume of the box is 200in^3 what are the dimentions of the sheet metal

OpenStudy (anonymous):

vol = 200 = 2 * x^2 x = 10 so side of original square = 10 + 2 = 12 ins

OpenStudy (anonymous):

the answer is 14x14 though

OpenStudy (anonymous):

oh yes - my mistake its 10+2+2 = 14 sorry

OpenStudy (anonymous):



OpenStudy (anonymous):

the above (not very good) drawing illustrates this

OpenStudy (anonymous):

welli still dont get how you figured it out :/